# Spin-operator form factors of the critical Ising chain and their finite volume scaling limits

### v3

@@ 2.2 The Wick contraction
To show this, we can change variable as
$s_{i}=\frac{t_{i}+1}{t_{i}-1},\ \frac{s_{i}-s_{j}}{s_{i}s_{j}-1}\rightarrow\frac{t_{i}-t_{j}}{t_{i}+t_{j}}\ ,$ (178)
and then the Eq. (177) becomes a Cauchy-type identity. Using Eq. (177), the contractions with the spin operator leads to an extra 1-1 sign, as well as a product of the $\frac{1}{\sqrt{z}}$s from the $NS$ sector, and a product of the $\sqrt{\beta}$s from the R sector. These factors compensate the overall minus sign, and the additional factors in $f_{\pm}$. As such, one obtains
@@ 2.3 The vacuum overlap
$e^{-\Delta_{E}(N)\tau}\langle\Omega_{R}^{+}|c_{1}\sigma(\tau)\sigma(0)c_{1}^{\dagger}|\Omega_{R}^{+}\rangle=|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}\det(1+K_{\tau})\ ,$ (182)
@@
Thus, it is now a ratio of two Cauchy determinants. Here we compute first the simple one,
@@
We already encountered these products before when showing the Eq. (63)
@@
We now compute the determinant. Since the $|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}$ is positive, here we will not care two much about the signs
@@
We now derive the integral representation Eq. (81) that allows to expand the overlap to an arbitrary order in $\frac{1}{N}$. We start from Eq. (75). Here the key observation is that for $\beta\in S_{+}$ and $\beta\neq 1$, as well as for $z\in S_{-}$, one always has ${\rm Re}(-i\sqrt{z}),{\rm Re}(-i\sqrt{\beta)}>0$. As such, it is convenient to introduce a Barnes representation in the variable $i\sqrt{\beta}$. Denote the Barnes parameter as $u$, one has for ${\cal Z}=z,\beta$
@@
Now, as far as ${\cal Z}\notin R_{+}$, the exponential decay in the ${\rm Im}(u)$ direction will be preserved and the double integral is absolutely convergent. The $t$ integral can be performed, leading to
@@
It is easy to perform the finite sums and obtain
@@
Integrals here are elementary and can be integrate to $\zeta$ values at positive even integers. As such, in the large $N$ limit, one has the following large $N$ expansion
@@
To show the product formula in Eq. (212), simplify notice that products of factorials can be expressed through hyper-factorials $H(n)=\prod_{i=1}^{n}i^{i}$, whose large $n$ asymptotics leads to the Glaisher constant
@@
$\int_{0<{\rm Re}(z)<1}\frac{dz}{2\pi i}(2-2^{z})(2-2^{-z})\Gamma(z)\Gamma(-z)\zeta(z)\zeta(-z)=\ln\bigg(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\bigg)\ .$ (215)
Notice that the integrand is symmetric under $z\rightarrow-z$. To be more precise, one first write
@@
$=\sum_{i=0}^{\infty}\bigg(\ln\sqrt{i+\frac{1}{2}}+\ln\frac{\Gamma\left(i+1\right)}{\Gamma(i+\frac{3}{2})}-\ln\sqrt{i+1}-\ln\frac{\Gamma\left(i+\frac{3}{2}\right)}{\Gamma(i+2)}\bigg)$
$=\ln\left(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\right)\ .$ (216)
@@
Finally, we use the functional equation of the $\zeta$ function to change $\zeta(u+1)$ to $\zeta(-u)$ and obtain Eq. (215). In Appendix. A, we will use this Barnes integral to find the small mass expansion of a related crossover function.
@@ 2.4 Scaling limit of the two-point correlator
As the $N\rightarrow\infty$, we have shown that the overlap between the two ground states approaches 0 at a very slow speed $N^{-\frac{1}{4}}$ controlled by the scaling dimension of the spin operator. The rescaled spin operator $\hat{\sigma}$ in Eq. (53), on the other hand, allows the finite volume scaling limit at the level of the form factors, given by the Eq. (107), Eq. (108) and Eq. (109). Here we would like to show that the finite-volume scaling limit at the level of the spin-spin two-point correlator also exists and leads to the standard CFT correlator on a cylinder. More precisely, we separate two spin operators by an Euclidean time like in Eq. (180), but between two NS ground states. Then, the finite volume scaling limit means
$\tau\rightarrow\infty\ ,N\rightarrow\infty,\ \frac{r}{L}=\frac{\tau}{N}={\cal O}(1)\ .$ (221)
We first show that the scaling limit of the form factor expansion can be taken under the summation for the individual form factors. For this, first notice that the exponential factors can be bounded uniformly by
@@
Second, notice that if $\beta,\beta^{\prime}\in S_{+}$, then
@@
$\sum_{k=0}^{\frac{N-1}{2}}\frac{1}{(2k+1)!}\sum_{\beta_{i}\in S_{+}}\prod_{i}e^{-\epsilon(\beta_{i})\tau}|d(\beta_{i})|^{2}\prod_{i<j}\left(\frac{\sqrt{\beta}_{i}-\sqrt{\beta_{j}}}{\sqrt{\beta_{i}}\sqrt{\beta_{j}}-1}\right)^{2}$
$\leq\sum_{k=0}^{\infty}\frac{1}{(2k+1)!}\left(c^{2}\sum^{\infty}_{k=-\infty}e^{-\frac{\pi r|k|}{L}}\right)^{2k+1}<\infty\ ,$ (225)
allowing to take the scaling limit under the sum, due to the dominated convergence theorem.
After showing the convergence, we show that the rescaled two point correlator can be expressed as the square of a Fredholm determinant. In fact, in the product formulas Eq. (107), Eq. (108), there are no correlations between the left and right momentum. As such, in the vacuum two point correlator, within each of the four possible even odd parity assignment of the excitation numbers, the particle numbers in different L, R sectors can be summed independently, leading to
@@
$F_{-}(w)=\sum_{k=0}^{\infty}\frac{1}{(2k+1)!}\sum_{\{n_{i}\geq 1\}}\prod_{i=1}^{2k+1}\tau^{2}(n_{i})w^{n_{i}}\prod_{i<j}\left(\frac{n_{i}-n_{j}}{n_{i}+n_{j}}\right)^{2}\ .$ (228)
Here, $\tau(n)$ is defined in Eq. (98), and we have introduced the notation
@@
Using the Cauchy determinant formula, it is nor hard to show that the $F_{+}+F_{-}$ can be combined to a single Fredholm determinant
@@
This is another major result of the work.
@@
This is exactly the standard CFT two-point correlator on an infinite Cylinder. As such, after multiplying back the factor $|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}$, at small $r\ll L$, one has then
@@
This is again consistent with the standard CFT formula Di Francesco et al. (1997).
@@ 3 Comparison with literature and further comments
where $\mu\neq 0$ is the fermion mass. At any fixed $\theta$, the $\mu\rightarrow 0$ limit diverges logarithmically. However, in order to reduce to the CFT limit, one must fix the quantization levels and the radius, then send $\mu$ to zero, which means $\theta\rightarrow\pm\infty$ and suppresses the $\theta^{\prime}={\cal O}(1)$ region. On the other hand, when $\theta^{\prime}-\theta={\cal O}(1)$, then the denominator is no longer large, and the $\mu$ in the exponential is also be compensated by the $\mu L\cosh\theta^{\prime}\sim\mu L\cosh\theta$, which is $O(1)$ in the scaling limit. To magnify the $\theta^{\prime}-\theta={\cal O}(1)$ region, we change the variable to the momentum space, and denote $p=\mu\sinh\theta,E=\mu\cosh\theta$
@@
The first line is essentially the integral given in Gabai and Yin (2022), while in the second line, we have averaged with the $k\rightarrow-k$ integrand and changed to $E^{\prime}=\sqrt{k^{2}+\mu^{2}}$. In this form, it finally becomes transparent that there is a non-trivial $\mu\rightarrow 0$ limit at fixed ELEL
@@
where $\tau$ is given by Eq. (98). The above exactly reproduces the dressing functions in Eq. (107) and Eq. (108). For the zero mode, instead one needs
$\kappa(\mu L,\mu L)=\frac{1}{\pi}\int_{1}^{\infty}\frac{dt}{t\sqrt{t^{2}-1}}\ln\frac{1-e^{-\mu Lt}}{1+e^{-\mu Lt}}=\frac{1}{2}\ln\mu L+{o}(1)\ ,$ (243)
implying
$\frac{e^{\kappa(0)}}{\sqrt{\mu L}}\rightarrow 1\ ,$ (244)
@@
Second, in the introduction, we have mentioned that in Gehlen et al. (2008), the authors have omitted the $g=1$ point in their discussions, but this does not rule out the possibility that the results therein also applies at the $g=1$ by taking limit. Let’s show that this is indeed the case. The major quantity to check are the products of their leg function. In our notation, the leg function in Gehlen et al. (2008) reads
@@
where we have used the relation p(1)q(1)=2ip(1)q(1)=-2i in Eq. (71). As such, all the form factor squared in Gehlen et al. (2008); Iorgov et al. (2011), in the proper $g\rightarrow 1^{+}$ limit approached from the disordered side, agree with the results derived directly at the critical point.
Third, we should comment that although we believe the Fredholm determinant Eq. (231) leads exactly to the CFT correlator through the formula
@@
we have not yet verified this analytically. However, we have noticed that the form factor expansion of this determinant is similar to the infinite-volume form factor expansion Lyberg and McCoy (2007) of the Ising chain or diagonal Ising correlators away from the critical point, and at zero distance. For example, the ${\rm Tr}{\cal K}$ equals to
@@
where $K(w)$ is the complete Elliptic integral of the second Kind. The $\frac{K(w)}{\pi}$ is exactly half of the one particle form factor contribution $f^{(1)}_{0,0}(t)$ to the diagonal correlator at zero separation Boukraa et al. (2007). If such relations could be established at high orders, then the identity Eq. (254) becomes understandable: $\sigma_{00}^{2}=1$, while the form factor expansion in Boukraa et al. (2007) contains an overall $(1-t)^{\frac{1}{4}}$. This implies that all the form factors at zero separation must sum to $(1-t)^{-\frac{1}{4}}$. It is also possible that differential equations in $w$ could be established, similar to the Painleve VI for the parameter derivatives Jimbo and Miwa (1980). Finally, it is also interesting to see how the form factors can be summed to the more general toroidal correlators at a finite temperature.

### v4

@@ 2.2 The Wick contraction
To show this, we can change variables as
$s_{i}=\frac{1+t_{i}}{1-t_{i}},\ \frac{s_{i}-s_{j}}{s_{i}s_{j}-1}\rightarrow\frac{t_{i}-t_{j}}{t_{i}+t_{j}}\ ,$ (182)
and then the Eq. (181) becomes a Cauchy-type identity. Using Eq. (181), the contractions with the spin operator leads to an extra 1-1 sign, as well as a product of all the $\frac{1}{\sqrt{z}}$s from the $NS$ sector, and another of all the $\sqrt{\beta}$s from the R sector. These factors compensate the overall minus sign, and the additional factors in $f_{\pm}$. As such, one obtains
@@ 2.3 The vacuum overlap
$e^{-\Delta_{E}(N)\tau}\langle\Omega_{R}|c_{1}\sigma(\tau)\sigma(0)c_{1}^{\dagger}|\Omega_{R}\rangle=|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}\det(1+K_{\tau})\ ,$ (186)
@@
Thus, it is now a ratio of two Cauchy determinants. We first consider the simple one, the inverse of which read
@@
We already encountered these unrestricted products before when showing the Eq. (65)
@@
From the inverse, the determinant can be easily extracted. Since the $|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}$ is positive, here we will not care too much about the signs
@@
We now derive the integral representation Eq. (83) that allows to expand the overlap to an arbitrary order in $\frac{1}{N}$. We start from Eq. (77). Here the key observation is that for $\beta\in S_{+}/\{1\}$, as well as for $z\in S_{-}$, one always has ${\rm Re}(-i\sqrt{z}),{\rm Re}(-i\sqrt{\beta)}>0$. As such, it is convenient to introduce a Barnes representation in the variable $i\sqrt{\beta}$. Denote the Barnes parameter as $u$, by performing the Mellin transform under the integral representations Eq. (78), for ${\cal Z}\notin R_{+}$ one has
@@
Now, as far as ${\cal Z}\notin R_{+}$, the exponential decay in the ${\rm Im}(u)$ direction will be preserved and the double integral is absolutely convergent. The Barnes representation also simplifies the $t$ dependencies, allowing the $t$ integral to be performed explicitly as
@@
It is easy to perform the finite sums which are geometric and obtain
@@
Integrals here are elementary and can be integrate to $\zeta$ values at positive even integers. As such, in the large $N$ limit, one has the following expansion
@@
To show the product formula in Eq. (216), simplify notice that products of factorials can be expressed through hyper-factorials $H(n)=\prod_{i=1}^{n}i^{i}$, whose large $n$ asymptotics lead to the Glaisher constant
@@
$\int_{c-i\infty}^{c+i\infty}\frac{du}{2\pi i}(2-2^{u})(2-2^{-u})\Gamma(u)\Gamma(-u)\zeta(u)\zeta(-u)=\ln\bigg(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\bigg)\ ,$ (219)
where $0<c<1$. Notice that the integrand is symmetric under $z\rightarrow-z$. To be more precise, one first write
@@
$=\sum_{i=0}^{\infty}\bigg(\ln\frac{\sqrt{i+\frac{1}{2}}}{\sqrt{i+1}}+\ln\frac{\Gamma\left(i+1\right)}{\Gamma(i+\frac{3}{2})}-\ln\frac{\Gamma\left(i+\frac{3}{2}\right)}{\Gamma(i+2)}\bigg)=\ln\left(\pi^{\frac{1}{4}}G\left(\frac{1}{2}\right)G\left(\frac{3}{2}\right)\right)\ .$ (220)
@@
Finally, we use the functional equation of the $\zeta$ function to change $\zeta(u+1)$ to $\zeta(-u)$ and obtain Eq. (219). In Appendix. A, as an application of this Barnes integral, we compute the small mass expansion of a related crossover function.
@@ 2.4 Scaling limit of the two-point correlator
As the $N\rightarrow\infty$, we have shown that the overlap between the two ground states approaches 0 at a very slow speed $N^{-\frac{1}{4}}$ controlled by the scaling dimension of the spin operator. The rescaled spin operator $\hat{\sigma}$ in Eq. (55), on the other hand, allows the finite volume scaling limit at the level of the form factors, given by the Eq. (109), Eq. (110) and Eq. (111). Here we would like to show that the finite-volume scaling limit at the level of the spin-spin two-point correlator also exists and leads to the standard CFT correlator on a cylinder. More precisely, we first separate the two spin operators by an Euclidean time like in Eq. (184), but within two NS ground states. Then, the finite volume scaling limit means
$\tau\rightarrow+\infty\ ,N\rightarrow+\infty,\ \frac{r}{L}=\frac{\tau}{N}={\cal O}(1)\ .$ (225)
We first show that the scaling limit of the form factor expansion can be taken under the summation for the individual form factors. For this, first notice that the time evolution factors can be bounded uniformly by
@@
Second, notice that if $\beta,\beta^{\prime}\in S_{+}/\{1\}$, then
@@
$e^{\Delta_{E}(N)\tau}\langle\Omega_{NS}|\hat{\sigma}(\tau)\hat{\sigma}(0)|\Omega_{NS}\rangle$
[…]
$+$ $1+\sum_{k=1}^{\frac{N-2}{2}}\frac{1}{(2k)!}\sum_{\beta_{i}\in S_{+}/\{1\}}\prod_{i=1}^{2k}e^{-\epsilon(\beta_{i})\tau}|d(\beta_{i})|^{2}\prod_{i<j}\left(\frac{\sqrt{\beta}_{i}-\sqrt{\beta_{j}}}{\sqrt{\beta_{i}}\sqrt{\beta_{j}}-1}\right)^{2}$
$\leq\sum_{k=0}^{\infty}\frac{1}{k!}\left(2c^{2}\sum^{\infty}_{n=1}e^{-\frac{\pi rn}{L}}\right)^{k}<\infty\ .$ (229)
Notice that the second line is the contribution from intermediate states without $\beta=1$ (zero momentum) excitations, while the third line is from states with a $\beta=1$ excitation. In particular, the +1+1 in the third line is due to the intermediate state $c_{1}^{\dagger}|\Omega_{R}\rangle$, and we have used the fact $d(1)=1$ in the third line. The bound Eq. (229) allows to take the scaling limit under the sum, due to the dominated convergence theorem.
After showing the convergence, we show that the rescaled two point correlator can be expressed as the square of a Fredholm determinant. In fact, in the product formulas Eq. (109), Eq. (110), there are no correlations between the left and right momentum. As such, in the vacuum two point correlator, within each of the four possible even odd parity assignments of the excitation numbers, the particle numbers in different L, R sectors can be summed independently, leading to
@@
$F_{-}(w)=\sum_{k=0}^{\infty}\frac{1}{(2k+1)!}\sum_{\{n_{i}\geq 1\}}\prod_{i=1}^{2k+1}\tau^{2}(n_{i})w^{n_{i}}\prod_{i<j}\left(\frac{n_{i}-n_{j}}{n_{i}+n_{j}}\right)^{2}\ ,$ (232)
where $\tau(n)$ is defined in Eq. (100), and we have introduced the notation
@@
In fact, the $2F_{+}F_{-}$ is the scaling limit for the second line of Eq. (229), while $F_{+}^{2}+F_{-}^{2}$ is for the third line. Using the Cauchy determinant formula, it is nor hard to show that the $F_{+}+F_{-}$ can be combined to a single Fredholm determinant
@@
This is another major result of the work. Notice that after the similarity transformation $iw^{j}\rightarrow\sqrt{ij}w^{\frac{i+j}{2}}$, the operator $\cal K$} is clearly of trace-class when $|w|<1$, and the Fredholm determinant is well defined.
@@
This is exactly the standard CFT two-point correlator on an infinite cylinder Di Francesco et al. (1988, 1987, 1997) and generalizes the sum-rules in Di Francesco et al. (1987). As such, after multiplying back the factor $|\langle\Omega_{R}|\Omega_{NS}\rangle|^{2}$, at small $r\ll L$, one has then
@@
This is again consistent with the standard CFT formula Di Francesco et al. (1988, 1997).
@@ 3 Comparison with literature and further comments
where $\mu>0$ is the fermion mass. At any fixed $\theta$, the $\mu\rightarrow 0$ limit diverges logarithmically. However, in order to reach the CFT limit, one must fix the quantization levels and the radius, then send $\mu$ to zero, which means $\theta\rightarrow\pm\infty$ and suppresses the $\theta^{\prime}={\cal O}(1)$ region. On the other hand, when $\theta^{\prime}-\theta={\cal O}(1)$, then the denominator is no longer large, and the $\mu$ in the exponential is also compensated by the $\mu L\cosh\theta^{\prime}\sim\mu L\cosh\theta$, which is $O(1)$ in the scaling limit. To magnify the $\theta^{\prime}-\theta={\cal O}(1)$ region, we change the variable to the momentum space, and denote $p=\mu\sinh\theta,E=\mu\cosh\theta$
@@
The first line is essentially the integral given in Gabai and Yin (2022), while in the second line, we have averaged with the $k\rightarrow-k$ integrand and changed to $E^{\prime}=\sqrt{k^{2}+\mu^{2}}$. In this form, it finally becomes transparent that there is a non-trivial $\mu\rightarrow 0^{+}$ limit at fixed ELEL
@@
where $\tau$ is given by Eq. (100). The above exactly reproduce the dressing functions in Eq. (109) and Eq. (110). For the zero mode, instead one needs
$\kappa(\mu L,\mu L)=\frac{1}{\pi}\int_{1}^{\infty}\frac{dt}{t\sqrt{t^{2}-1}}\ln\frac{1-e^{-\mu Lt}}{1+e^{-\mu Lt}}=\frac{1}{2}\ln\mu L+{o}(1)\ ,\ \frac{e^{\kappa(0)}}{\sqrt{\mu L}}\rightarrow 1\ ,$ (247)
@@
Second, in the introduction, we have mentioned that in Gehlen et al. (2008), the authors have omitted the $g=1$ point in their discussions, but this does not rule out the possibility that the results therein also apply at the $g=1$ point by taking limit. Let’s show that this is indeed the case. The major quantity to check are the products of their leg function. In our notation, the leg function in Gehlen et al. (2008) for periodic momenta reads
@@
where we have used the relation p(1)q(1)=2ip(1)q(1)=-2i in Eq. (73). As such, all the form factors squared in Gehlen et al. (2008); Iorgov et al. (2011), in the proper $g\rightarrow 1^{+}$ limit approached from the disordered side, agree with the results derived directly at the critical point.
After comparing with the literature, we would like to make a few more comments on possible future directions and generalizations.
  1. 1.
    First, we should emphasize that although we believe that the Fredholm determinant Eq. (235) leads exactly to the CFT correlator through the formula
@@
    we have not yet verified this analytically. However, we have noticed that the expansion of this determinant is similar to the infinite-volume form factor expansion Lyberg and McCoy (2007) of the Ising chain or diagonal Ising correlators away from the critical point, and at zero or very small distances. For example, the first two terms in the expansion equal to
@@
    where $K(w)$ is the complete Elliptic integral of the second Kind. The $\frac{K(w)}{\pi}$ is exactly half of the one particle form factor contribution $f^{(1)}_{0,0}(t)$ to the diagonal correlator at zero separation Boukraa et al. (2007). If such relations could be established at high orders, then the identity Eq. (257) becomes understandable: $\sigma_{00}^{2}=1$, while the form factor expansion in Boukraa et al. (2007) contains an overall $(1-t)^{\frac{1}{4}}$. This implies that all the form factors at zero separation must sum to $(1-t)^{-\frac{1}{4}}$. It is also possible that differential equations in $w$ could be established, similar to the Painleve VI for the parameter derivatives Jimbo and Miwa (1980).
  2. 2.
    Second, although in this work we only checked the two-point correlator, given the finite-volume form-factors, it should be possible to show the convergence of the scaling limit and establish determinantal representations for multi-point spin-correlators at non-vanishing Euclidean time separations, in a way similar to the massive version in the infinite volume McCoy et al. (1977); Palmer and Tracy (1981). What remains not clear at the moment, is how to use the form factor representations, to show the resulting correlators indeed agree with the known CFT results.
    On the other hand, it was proven recently that the spin correlators of the 2D critical Ising model on a torus indeed converge to the CFT version Bayraktaroglu and Izyurov (2025), using more subtle methods that are not based on form factors. Meanwhile, finite-volume form factors of the critical Ising model can still be established using the strategy of Abraham (1978a, b) or in the formalism of Iorgov and Lisovyy (2011) based on elliptic Cauchy determinants, by noticing similar near Cauchy property at the critical point [^6] The author thanks Oleg Lisovyi for pointing this out in private communications.. They are expected to converge to the same set of CFT matrix elements, leading to identical form factor representations of rescaled correlators as obtained from the Ising chain. As such, by comparing the form factor representations and the known CFT results allows to establish many Fredholm determinant identities that are not obvious at the first glance at all. Clearly, it is also interesting to see how the form factors can be summed to the more general toroidal correlators at a finite temperature.
  3. 3.
    Third, we comment that the finite volume form-factors are also known explicitly for more general $XY$ spin-chains in a transverse field Iorgov (2011) away from the critical points. Either by taking the appropriate $XX$ limit ($\kappa=h=0$) or by deriving the finite-volume form factors directly at that point which is technically possible, one should be able to obtain form factors of certain vertex operators in the $c=1,\ \beta^{2}=4\pi$ Sine-Gordon CFT in the massless Dirac fermion basis.
  4. 4.
    Finally, we should mention that for integrable lattice models solved by the Algebraic Bethe-Ansartz (ABA) such as the $XXZ$ spin-chains, there are also discussions on the large distance properties based on form-factors Dugave et al. (2014); Göhmann et al. (2020a, b); Kozlowski (2019); Göhmann and Kozlowski (2024). These discussions mostly concern the infinite-volume limit at zero or non-zero temperatures, but the region $1\leq\Delta<1$ with CFT-like asymptotics are covered.
    It is interesting to see how to extract from the $XXZ$ spin-chain, the CFT matrix elements in more general $c=1$ Sine-Gordon CFTs in the ABA-type massless basis, by starting with a finite volume, then taking the finite-volume scaling limit. It could be expected that going to a finite volume would simplify a bit the analysis aspects (taking limits) of the massless form factors. Such matrix elements might also be related to the finite temperature results in Babenko et al. (2021); Göhmann and Kozlowski (2024); Göhmann et al. (2023). Showing the convergence to the cylindrical CFT correlators is expected to be more challenging than to obtain the form factors.